# Boaz Moerman

### v1

@@ 2.3. Divisors on toric pairs
This is the restriction of the pullback homomorphism in [Moe25conjecture, Definition 4.15] to torus-invariant divisors. Using the pullback, we define the Picard group of a toric pair.
@@
In [Moe25conjecture], we defined the Picard group and the effective cone more generally by using the divisor group
@@
By [CLS11, Theorem 4.1.3] every divisor on $X$ is linearly equivalent to a torus-invariant divisor. This directly implies that every divisor on $(X,M)$ is linearly equivalent to a torus-invariant divisor on $(X,M)$.
The proof of the statement for effective divisors is based on the proof of the analogous statements for divisors on $X$ given in [CLS11, Proposition 4.3.2, Lemma 15.1.8]. For an effective divisor $D$ on $(X,M)$, let $D^{\prime}$ be a torus-invariant divisor linearly equivalent to it. Let
@@
where we write $E\geq 0$ for a divisor $E$ on $(X,M)$ to indicate that the divisor is effective. Then $H^{0}(X,D^{\prime})$ is a vector space invariant under the natural torus action on $K(X)^{\times}$. Hence by [CLS11, Lemma 1.1.16],
@@
Now we will give a simple presentation for the Picard group of a toric pair, for this we first recall the homomorphism $\phi$ from [Moe24, Definition 5.10].
Definition 2.13.
Let $X$ be a smooth split toric variety with torus invariant prime divisors $D_{1},\dots,D_{n}$ and corresponding ray generators $\rho_{1},\dots,\rho_{n}$. We define $\phi\colon\mathbb{Z}^{n}\rightarrow N$ to be the group homomorphism given by
$(m_{1},\dots,m_{n})\rightarrow\sum_{i=1}^{n}m_{i}\rho_{i}.$
Using this map, we can compute the Picard group can be calculated using torus-invariant divisors using an analogue of the exact sequence 2.1.
@@
The torus-invariant principal divisors on $(X,M)$ are exactly the pullbacks of torus-invariant principal divisors on $X$. Thus, the kernel of this homomorphism is the image of $N^{\vee}$ in $\operatorname{Div}_{T}(X,M)$ by [CLS11, Theorem 4.1.3].
@@
In [Moe25conjecture, Definition 4.33], we introduced the canonical divisor class $K_{(X,M)}\in\operatorname{Pic}(X,M)$ of a pair. On toric pairs, this class has a particularly nice representative.
@@
By [CLS11, Theorem 8.2.3], the canonical divisor class of a toric variety is
@@
We recall the Fujita invariant and the $b$-invariant from [Moe25conjecture].
@@
Note that there need not exist a $\mathbb{Q}$-divisor class $L$ such that $\operatorname{pr}^{*}L=-K_{(X,M)}$. In particular, the $b$-invariant can be strictly smaller than the Picard rank of $(X,M)$ for all choices of $L$. For instance this happens for pairs corresponding to Campana points, see [Moe25conjecture, Example Example 4.38]
In [Moe25conjecture], rigid divisors on pairs were defined. An effective divisor on a pair $(X,M)$ is rigid if it is the only effective divisor in its linear equivalence class [Moe25conjecture, Definition 4.44]. In the proof of Theorem 3.3, we will consider rigid divisor as well as well as divisors which are rigid with respect to torus-invariant divisors.
@@
The pullback map is injective on a proper toric pair by [Moe25conjecture, Proposition 4.30.]. Thus, on such a pair, toric rigid divisors are the same as rigid divisors.
@@
Let $X={\mathbf{Proj}}_{\mathbb{P}^{1}}(\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(d))$ be the Hirzebruch surface of degree $d>0$. Let $D_{2}$ be the unique prime divisor with self-intersection $D_{2}^{2}=-d$, and let $D_{1},D_{3}$ be the torus-invariant prime divisors intersecting the prime divisor $D_{2}$. Let $(X,M)$ be the toric pair corresponding to the open $U^{\prime}=X\setminus(D_{1}\cup D_{3})$. We identify $\operatorname{Pic}(X,M)$ with $\operatorname{Pic}(U^{\prime})$ using the natural isomorphism as in [Moe25conjecture, Example 4.26]. Under this identification, the anticanonical divisor on $(X,M)$ is $K_{(X,M)}=[D_{2}]+[D_{4}]\in\operatorname{Pic}(U^{\prime})$, which is a nonzero effective divisor class. Since $\operatorname{Pic}(U^{\prime})=\mathbb{Z}$, every divisor on $U^{\prime}$ is linearly equivalent to a rational multiple of $K_{(X,M)}$. Consequently, any big and nef $\mathbb{Q}$-divisor on $X$ is toric adjoint rigid with respect to $(X,M)$.
@@ 2.4. Quasi-proper pairs
[…]
@@
which corresponds to $\sum_{i=1}^{n}\left(a+1-\frac{1}{m}\right)D_{i}$ under the identification in [Moe25conjecture, Example 4.25]. This implies $a((X,M),L)=a((X,\tilde{M}),L)=1$, as desired.
@@
As a special case of the previous example, we can take $X=\operatorname{Bl}_{(0:0:1)}\mathbb{P}^{2}$, let $D_{1}$ be the exceptional divisor and take $\mathfrak{M}=\{0\}$. Then the $\mathbb{Q}$-divisor class $L=K_{X}+aD$ is big and nef for any $a\in(-1,1]\cap\mathbb{Q}$, and $(X,M)$ is quasi-proper with respect to $L$.
@@ 2.5. Fans and universal torsors for pairs
In [Sal98, Section 11], Salberger verifies Manin’s conjecture for split toric varieties with the anticanonical height. In order to achieve this, he reduces the counting problem to estimating the volume of a domain $D(B)\subset Y(\mathbb{R})$ in the real locus of the universal torsor $Y$ of the variety. Afterwards, he uses the fan $\Sigma$ of $X$ to give a partition of $D(B)$ into simpler pieces: $D(B)=\bigcup_{\sigma\in\Sigma_{\max}}D(B,\sigma)$. In order to adapt the proof of Salberger to toric pairs, we thus need appropriate analogues of the fan and of the universal torsor for toric pairs.
@@
We call $Y_{M}$ the universal torsor of $(X,M)$, as it plays an analogous role to the universal torsor of $X$ in Salberger’s work [Sal98].
@@ 3.1. Heights
Let $X$ be a smooth proper split toric variety over $\mathbb{Q}$ with fan $\Sigma$. Any $\mathbb{Q}$-divisor class $L$ on $X$ naturally induces a height function $H_{L}$ on the torus $U\subset X$, as defined for divisor classes by Batyrev and Tschinkel in [BaTs95, Definition 2.1.7]. In this section, we describe this height using the description of this height given in [PiSc24, Section 6.3].
@@
The first part follows directly from the definition of $l^{(i)}(\mathbf{e}_{j})$. If $L$ is nef, then it is semiample by [CLS11, Theorem 6.3.12.], and the nonnegativity of $l^{(i)}(\mathbf{e}_{j})\geq 0$ is proved as in [Sal98, Proposition 8.7(a)]. ∎
We will now assume that $L$ is big and nef. As in [PiSc24, Section 6], we define the function
@@
for every $j=1,\dots,k$ and $(x_{1},\dots,x_{n})\in Y(\mathbb{Q})$. Now [PiSc24, Proposition 6.10] implies that the height of a point $(x_{1}\colon\dots\colon x_{n})\in X(\mathbb{Q})$ with respect to $L$, as defined in [PiSc24, Section 6.3], is given as
@@ 3.2. General asymptotics and the leading constant
In this section we present Theorem 3.3, which will describe the asymptotic behavior of the toric counting function $N_{(X,M),L,S}(B)$ on a quasi-proper pair. To describe the leading constant appearing in the asymptotic, we first recall the $\alpha$-constant of the pair relative to $L$ as defined in [Moe25conjecture, Definiton 8.2].
@@
Now we will state the main result of this article, which gives an asymptotic formula for the number of $\mathcal{M}$-points of bounded height on a quasi-proper toric pair $(X,M)$. If $L$ is adjoint rigid, then we furthermore give a simple description of the leading constant. In Theorem 3.7, we give a geometric interpretation of this constant, which shows that it is in agreement with [Moe25conjecture, Conjecture 1.4]. Similarly, we also describe the leading constant if $L$ is toric adjoint rigid.
@@
where $\Sigma_{\overline{M^{\circ}},\max}$ is the collection of maximal cones in the fan $\Sigma_{\overline{M^{\circ}}}$ introduced in Notation 2.27, $I(\sigma)$ is the index of the subgroup $\langle[\tilde{D}_{\mathbf{m}}]\mid\phi(\mathbf{m})\not\in\sigma\rangle$ in $\operatorname{Pic}(X,M^{\circ}(\sigma))$ and
@@
The proof of Theorem 3.3 is based on the proofs of Salberger [Sal98] and de la Bretèche [Bre01Toric] for Manin’s conjecture for split toric varieties with the anticanonical height.
@@ 3.3. Examples
In [PSTVA21, Section 3.2.1], the weak Campana points on $(\mathbb{P}^{2},\frac{1}{2}D_{1}+\frac{1}{2}D_{2}+\frac{1}{2}D_{3})$ over $\mathbb{Z}$ are considered and compared to the Campana points on the same Campana pair. The set of weak Campana points is
@@
while the set of Campana points is the subset consisting of the points $(x:y:z)$ such that each integer $x,y,z$ is squareful. In [PSTVA21, Propostition 3.6] it is proven that the number of weak Campana points with Weil height at most $B$ and $xyz\neq 0$ is at least $c_{1}B^{3/2}\log B$ as $B\rightarrow\infty$, for some constant $c_{1}>0$. Using Theorem 3.3, we can compute a precise asymptotic for the number of weak Campana points of bounded height. Let $(\mathbb{P}^{2}_{\mathbb{Q}},M)$ be the pair corresponding to the weak Campana points on $(\mathbb{P}^{2}_{\mathbb{Q}},\frac{1}{2}D_{1}+\frac{1}{2}D_{2}+\frac{1}{2}D_{3})$. For the divisor class $L=[D_{1}]\in\operatorname{Pic}(\mathbb{P}^{2}_{\mathbb{Q}})$, [Moe25conjecture, Lemma 4.39] implies that the pair $(\mathbb{P}^{2}_{\mathbb{Q}},M^{\circ})$ as in Definition 3.2 is contained in the proper pair $(X,M^{\prime})$ given by $\mathfrak{M}^{\prime}=\{(0,0,0),(2,0,0),(0,2,0),(0,0,2),(1,1,0),(1,0,1),(0,1,1)\}$. Since $\Gamma_{M^{\prime}}=\mathfrak{M}^{\prime}\setminus\{(0,0,0)\}$ and
@@
the adjoint divisor of $L$ with respect to $(X,M)$ is 0 so we must have $(\mathbb{P}^{2}_{\mathbb{Q}},M^{\circ})=(\mathbb{P}^{2}_{\mathbb{Q}},M^{\prime})$. Furthermore, this description of $\operatorname{pr}_{M^{\prime}}^{*}[D_{1}]$ also implies that $a((\mathbb{P}^{2}_{\mathbb{Q}},M),[D_{1}])=\frac{3}{2}$ and $b(\mathbb{Q},(X,M),[D_{1}])=\operatorname{rank}\operatorname{Pic}(X,M^{\circ})$. Since $D_{1}$, $D_{2}$ and $D_{3}$ are linearly equivalent to each other, the divisor $\tilde{D}_{(1,0,1)}$ is linearly equivalent to $2\tilde{D}_{(0,2,0)}-2\tilde{D}_{(0,0,2)}+\tilde{D}_{(1,1,0)}$ and similarly $\tilde{D}_{(0,1,1)}$ is linearly equivalent to $2\tilde{D}_{(2,0,0)}-2\tilde{D}_{(0,0,2)}+\tilde{D}_{(1,1,0)}$. By Proposition 2.14 these are the only relations between torus-invariant prime divisors on $(X,M^{\circ})$, so $\operatorname{Pic}(X,M^{\circ})\cong\mathbb{Z}^{4}$ is freely generated by $\tilde{D}_{(2,0,0)}$, $\tilde{D}_{(0,2,0)}$, $\tilde{D}_{(0,0,2)}$ and $\tilde{D}_{(1,1,0)}$, and thus $b(\mathbb{Q},(X,M),L)=4$. The effective cone of $(\mathbb{P}^{2}_{\mathbb{Q}},M^{\circ})$ is generated by the divisors $\tilde{D}_{(2,0,0)},\tilde{D}_{(0,2,0)},\tilde{D}_{(0,0,2)},\tilde{D}_{(1,1,0)},\tilde{D}_{(1,0,1)},\tilde{D}_{(0,1,1)}$ by Proposition 2.12. By subdividing the dual of the effective cone $\operatorname{Eff}^{1}(X,M)$ into two simplicial cones, we compute the $\alpha$-constant using [Bar93, Example 2.1]:
$\frac{\alpha((\mathbb{P}^{2}_{\mathbb{Q}},M),[D_{1}])}{a((\mathbb{P}^{2}_{\mathbb{Q}},M),[D_{1}])(b(\mathbb{Q},(X,M),[D_{1}])-1)!}=\frac{1}{48}.$
Since $a((\mathbb{P}^{2}_{\mathbb{Q}},M))D_{1}$ is linearly equivalent to $\frac{1}{2}\left(D_{1}+D_{2}+D_{3}\right)$, we find $C_{\infty}=4\cdot 4\cdot 3=48$.
@@
Using the description of the $b$-invariant for weak Campana points given in [Moe25conjecture, Theorem 1.7], we can also determine the asymptotic growth for the number of points on projective space for which the product of the coordinates is an $m$-full number, generalizing Example 3.4.
@@
Note that $N(B)$ is the number of weak Campana points on the Campana pair $\left(\mathbb{P}^{n-1}_{\mathbb{Q}},\sum_{i=1}^{n}\left(1-\frac{1}{m}\right)D_{i}\right)$ of height at most $B$, where the divisors $D_{1},\dots,D_{n}$ are the coordinate hyperplanes. In particular, the log-anticanonical divisor class is given by $\sum_{i=1}^{n}\frac{1}{m}[D_{i}]$ and [Moe25conjecture, Theorem 1.7] implies that the Fujita invariant is $n/m$, while the $b$-invariant is equal to

### v2

@@ 2.4. Divisors on toric pairs
This is the restriction of the pullback homomorphism in [24, Definition 4.15] to torus-invariant divisors. Using the pullback, we define the Picard group of a toric pair.
@@
In [24], we defined the Picard group and the effective cone more generally by using the divisor group
@@
By [12, Theorem 4.1.3] every divisor on $X$ is linearly equivalent to a torus-invariant divisor. This directly implies that every divisor on $(X,M)$ is linearly equivalent to a torus-invariant divisor on $(X,M)$.
The proof of the statement for effective divisors is based on the proof of the analogous statements for divisors on $X$ given in [12, Proposition 4.3.2, Lemma 15.1.8]. For an effective divisor $D$ on $(X,M)$, let $D^{\prime}$ be a torus-invariant divisor linearly equivalent to it. Let
@@
where we write $E\geq 0$ for a divisor $E$ on $(X,M)$ to indicate that the divisor is effective. Then $H^{0}(X,D^{\prime})$ is a vector space invariant under the natural torus action on $K(X)^{\times}$. Hence by [12, Lemma 1.1.16],
@@
Using the map $\phi\colon\mathbb{Z}^{n}\rightarrow N$ introduced in Definition 2.8 and an analogue of the exact sequence (2.1), we will give a simple presentation for the Picard group of a toric pair.
@@
The torus-invariant principal divisors on $(X,M)$ are exactly the pullbacks of torus-invariant principal divisors on $X$. Thus, the kernel of this homomorphism is the image of $N^{\vee}$ in $\operatorname{Div}_{T}(X,M)$ by [12, Theorem 4.1.3].
@@
In [24, Definition 4.33], we introduced the canonical divisor class $K_{(X,M)}\in\operatorname{Pic}(X,M)$ of a pair. On toric pairs, this class has a particularly nice representative.
@@
By [12, Theorem 8.2.3], the canonical divisor class of a toric variety is
@@
We recall the Fujita invariant and the $b$-invariant from [24].
@@
Note that there need not exist a $\mathbb{Q}$-divisor class $L$ such that $\operatorname{pr}^{*}L=-K_{(X,M)}$. In particular, the $b$-invariant can be strictly smaller than the Picard rank of $(X,M)$ for all choices of $L$. For instance this happens for pairs corresponding to Campana points, see [24, Example 4.38]
In [24], rigid divisors on pairs were defined. An effective divisor on a pair $(X,M)$ is rigid if it is the only effective divisor in its linear equivalence class [24, Definition 4.44]. In the proof of Theorem 3.3, we will consider rigid divisors as well as divisors which are rigid with respect to torus-invariant divisors.
@@
The pullback map is injective on a proper toric pair by [24, Proposition 4.30.]. Thus, on such a pair, toric rigid divisors are the same as rigid divisors.
@@
Let $X={\mathbf{Proj}}_{\mathbb{P}^{1}}(\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(d))$ be the Hirzebruch surface of degree $d>0$. Let $D_{2}$ be the unique prime divisor with self-intersection $D_{2}^{2}=-d$, and let $D_{1},D_{3}$ be the torus-invariant prime divisors intersecting the prime divisor $D_{2}$. Let $(X,M)$ be the toric pair corresponding to the open $U^{\prime}=X\setminus(D_{1}\cup D_{3})$. We identify $\operatorname{Pic}(X,M)$ with $\operatorname{Pic}(U^{\prime})$ using the natural isomorphism as in [24, Example 4.26]. Under this identification, the anticanonical divisor on $(X,M)$ is $K_{(X,M)}=[D_{2}]+[D_{4}]\in\operatorname{Pic}(U^{\prime})$, which is a nonzero effective divisor class. Since $\operatorname{Pic}(U^{\prime})=\mathbb{Z}$, every divisor on $U^{\prime}$ is linearly equivalent to a rational multiple of $K_{(X,M)}$. Consequently, any big and nef $\mathbb{Q}$-divisor on $X$ is toric adjoint rigid with respect to $(X,M)$.
@@ 2.5. Quasi-proper pairs
Let $X$ be a smooth proper variety and let $L\in\operatorname{Pic}(X)_{\mathbb{Q}}$ be a big and nef $\mathbb{Q}$-divisor class. A smooth pair $(X,M)$ is quasi-proper with respect to $L$ if there exists a smooth proper pair $(X,\tilde{M})\supset(X,M)$ such that $a((X,M),L)=a((X,\tilde{M}),L)$. Here the containment is as in [24, Definition 3.12] (which for toric pairs boils down to $\mathfrak{M}\subset\tilde{\mathfrak{M}}$).
@@
$\tilde{\mathfrak{M}}=\mathfrak{M}\cup\{m\mathbf{e}_{1},\dots,m\mathbf{e}_{n}\}$
for any sufficiently large integer $m$.
@@
which corresponds to $\sum_{i=1}^{n}\left(a+1-\frac{1}{m}\right)D_{i}$ under the identification in [24, Example 4.25]. This implies $a((X,M),L)=a((X,\tilde{M}),L)=1$, as desired.
@@
As a special case of the previous example, we can take $X=\operatorname{Bl}_{(0:0:1)}\mathbb{P}^{2}$, let $D_{1}$ be the exceptional divisor and take $\mathfrak{M}=\{0\}$. If $a\in(-1,1]\cap\mathbb{Q}$, then the $\mathbb{Q}$-divisor class $L=K_{X}+aD_{1}$ is big and nef, while $(X,M)$ is quasi-proper with respect to $L$.
@@ 2.6. Fans and universal torsors for pairs
In [30, Section 11], Salberger verifies Manin’s conjecture for split toric varieties with the anticanonical height. In order to achieve this, he reduces the counting problem to estimating the volume of a domain $D(B)\subset Y(\mathbb{R})$ in the real locus of the universal torsor $Y$ of the variety. Afterwards, he uses the fan $\Sigma$ of $X$ to give a partition of $D(B)$ into simpler pieces: $D(B)=\bigcup_{\sigma\in\Sigma_{\max}}D(B,\sigma)$. In order to adapt the proof of Salberger to toric pairs, we thus need appropriate analogues of the fan and of the universal torsor for toric pairs.
@@
We call $Y_{M}$ the universal torsor of $(X,M)$, as it plays an analogous role to the universal torsor of $X$ in Salberger’s work [30].
@@ 3.1. Heights
Let $X$ be a smooth proper split toric variety over $\mathbb{Q}$ with fan $\Sigma$. Any $\mathbb{Q}$-divisor class $L$ on $X$ naturally induces a height function $H_{L}$ on the torus $U\subset X$, as defined for divisor classes by Batyrev and Tschinkel in [5, Definition 2.1.7]. In this section, we describe this height using its description given in [28, Section 6.3].
@@
The first part follows directly from the definition of $l^{(i)}(\mathbf{e}_{j})$. If $L$ is nef, then it is semiample by [12, Theorem 6.3.12.], and the nonnegativity of $l^{(i)}(\mathbf{e}_{j})\geq 0$ is proved as in [30, Proposition 8.7(a)]. ∎
We will now assume that $L$ is big and nef. As in [28, Section 6], we define the function
@@
for every $j=1,\dots,k$ and $(x_{1},\dots,x_{n})\in Y(\mathbb{Q})$. Now [28, Proposition 6.10] implies that the height of a point $(x_{1}\colon\dots\colon x_{n})\in X(\mathbb{Q})$ with respect to $L$, as defined in [28, Section 6.3], is given as
@@ 3.2. General asymptotics and the leading constant
In this section we present Theorem 3.3, which will describe the asymptotic behavior of the toric counting function $N_{(X,M),L,S}(B)$ on a quasi-proper pair. To describe the leading constant appearing in the asymptotic, we first recall the $\alpha$-constant of the pair relative to $L$ as defined in [24, Definiton 8.2].
@@
Now we will state the main result of this article, which gives an asymptotic formula for the number of $\mathcal{M}$-points of bounded height on a quasi-proper toric pair $(X,M)$. If $L$ is adjoint rigid, then we furthermore give a simple description of the leading constant. In Theorem 3.8, we give a geometric interpretation of this constant, which shows that it is in agreement with [24, Conjecture 1.4]. Similarly, we also describe the leading constant if $L$ is toric adjoint rigid.
@@
where $\Sigma_{\overline{M^{\circ}},\max}$ is the collection of maximal cones in the fan $\Sigma_{\overline{M^{\circ}}}$ introduced in Notation 2.31, $I(\sigma)$ is the index of the subgroup $\langle[\tilde{D}_{\mathbf{m}}]\mid\phi(\mathbf{m})\not\in\sigma\rangle$ in $\operatorname{Pic}(X,M^{\circ}(\sigma))$ and
@@
The proof of Theorem 3.3 is based on the proofs of Salberger [30] and de la Bretèche [13] for Manin’s conjecture for split toric varieties with the anticanonical height.
@@ 3.3. Examples
Let $f\colon\mathbb{P}^{2}\dashrightarrow\mathbb{P}^{2}$ be the rational map given by $(x:y:z)\mapsto(x^{2}:y^{2}:xz)$ as in Example 1.6. We will derive the given asymptotic formula for the number of rational points of bounded height in the image. If $(\mathbb{P}^{2},M)$ is the toric pair given by $\mathfrak{M}=\mathbb{N}^{2}$, then we compute that $f_{*}\mathfrak{M}$ is the monoid generated by $(2,0,0),(0,2,0),(0,0,1),(1,1,0)$ and thus $\Gamma_{f_{*}M}=\{(2,0,0),(0,2,0),(0,0,1),(1,1,0)\}$. The Picard group $\operatorname{Pic}(X,f_{*}M)\cong\mathbb{Z}^{2}\oplus\mathbb{Z}/2\mathbb{Z}$ is generated by $\tilde{D}_{(2,0,0)},\tilde{D}_{(1,1,0)}$ together with the torsion element $\tilde{D}_{(2,0,0)}-\tilde{D}_{(0,2,0)}$. Furthermore, we have that
$D_{(\mathbb{P}^{2},f_{*}M)}=\operatorname{pr}^{*}_{f_{*}M}\left(\tfrac{1}{2}D_{1}+\tfrac{1}{2}D_{2}+D_{3}\right),$
so $a((\mathbb{P}^{2},f_{*}M),[D_{1}])=2$ and $b(\mathbb{Q},(\mathbb{P}^{2},f_{*}M),[D_{1}])=2$. Now Theorem 3.3 shows that the number of $f_{*}\mathcal{M}$-points of height at most $B$ tends to $B\left(Q(\log B)+O\left(B^{-\theta}\right)\right)$, where $Q$ is a linear polynomial with leading coefficient
$C=\frac{1/4}{2}\cdot 8\cdot\prod_{p\,\mathrm{prime}}(1-2p^{-2}+2p^{-3}-p^{-4}).$
Now the asymptotic stated in Example 1.6 follows by combining this result with Corollary 1.5, where we have $\Delta=2$ as $f$ has degree $2$.
Example 3.5.
In [29, Section 3.2.1], the weak Campana points on $(\mathbb{P}^{2},\frac{1}{2}D_{1}+\frac{1}{2}D_{2}+\frac{1}{2}D_{3})$ over $\mathbb{Z}$ are considered and compared to the Campana points on the same Campana pair. The set of weak Campana points is
@@
while the set of Campana points is the subset consisting of the points $(x:y:z)$ such that each integer $x,y,z$ is squareful. In [29, Propostition 3.6] it is proven that the number of weak Campana points with Weil height at most $B$ and $xyz\neq 0$ is at least $c_{1}B^{3/2}\log B$ as $B\rightarrow\infty$, for some constant $c_{1}>0$. Using Theorem 3.3, we can compute a precise asymptotic for the number of weak Campana points of bounded height. Let $(\mathbb{P}^{2},M)$ be the pair corresponding to the weak Campana points on $(\mathbb{P}^{2},\frac{1}{2}D_{1}+\frac{1}{2}D_{2}+\frac{1}{2}D_{3})$. For the divisor class $L=[D_{1}]\in\operatorname{Pic}(\mathbb{P}^{2})$, [24, Lemma 4.39] implies that the pair $(\mathbb{P}^{2},M^{\circ})$ as in Definition 3.2 is contained in the proper pair $(X,M^{\prime})$ given by $\mathfrak{M}^{\prime}=\{(0,0,0),(2,0,0),(0,2,0),(0,0,2),(1,1,0),(1,0,1),(0,1,1)\}$. Since $\Gamma_{M^{\prime}}=\mathfrak{M}^{\prime}\setminus\{(0,0,0)\}$ and
@@
the adjoint divisor of $L$ with respect to $(X,M)$ is 0 so we must have $(\mathbb{P}^{2},M^{\circ})=(\mathbb{P}^{2},M^{\prime})$. Furthermore, this description of $\operatorname{pr}_{M^{\prime}}^{*}[D_{1}]$ also implies that $a((\mathbb{P}^{2},M),[D_{1}])=\frac{3}{2}$ and $b(\mathbb{Q},(X,M),[D_{1}])=\operatorname{rank}\operatorname{Pic}(X,M^{\circ})$. Since $D_{1}$, $D_{2}$ and $D_{3}$ are linearly equivalent to each other, the divisor $\tilde{D}_{(1,0,1)}$ is linearly equivalent to $2\tilde{D}_{(0,2,0)}-2\tilde{D}_{(0,0,2)}+\tilde{D}_{(1,1,0)}$ and similarly $\tilde{D}_{(0,1,1)}$ is linearly equivalent to $2\tilde{D}_{(2,0,0)}-2\tilde{D}_{(0,0,2)}+\tilde{D}_{(1,1,0)}$. By Proposition 2.18 these are the only relations between torus-invariant prime divisors on $(X,M^{\circ})$, so $\operatorname{Pic}(X,M^{\circ})\cong\mathbb{Z}^{4}$ is freely generated by $\tilde{D}_{(2,0,0)}$, $\tilde{D}_{(0,2,0)}$, $\tilde{D}_{(0,0,2)}$ and $\tilde{D}_{(1,1,0)}$, and thus $b(\mathbb{Q},(X,M),L)=4$. The effective cone of $(\mathbb{P}^{2},M^{\circ})$ is generated by the divisors $\tilde{D}_{(2,0,0)},\tilde{D}_{(0,2,0)},\tilde{D}_{(0,0,2)},\tilde{D}_{(1,1,0)},\tilde{D}_{(1,0,1)},\tilde{D}_{(0,1,1)}$ by Proposition 2.17. By subdividing the dual of the effective cone $\operatorname{Eff}^{1}(X,M)$ into two simplicial cones, we compute the $\alpha$-constant using [3, Example 2.1]:
$\frac{\alpha((\mathbb{P}^{2},M),[D_{1}])}{a((\mathbb{P}^{2},M),[D_{1}])(b(\mathbb{Q},(X,M),[D_{1}])-1)!}=\frac{1}{48}.$
Since $a((\mathbb{P}^{2},M))D_{1}$ is linearly equivalent to $\frac{1}{2}\left(D_{1}+D_{2}+D_{3}\right)$, we find $C_{\infty}=4\cdot 4\cdot 3=48$.
@@
Using the description of the $b$-invariant for weak Campana points given in [24, Theorem 1.7], we can also determine the asymptotic growth for the number of points on projective space for which the product of the coordinates is an $m$-full number, generalizing Example 3.5.
@@
Note that $N(B)$ is the number of weak Campana points on the Campana pair $\left(\mathbb{P}^{n-1},\sum_{i=1}^{n}\left(1-\frac{1}{m}\right)D_{i}\right)$ of height at most $B$, where the divisors $D_{1},\dots,D_{n}$ are the coordinate hyperplanes. In particular, the log-anticanonical divisor class is given by $\sum_{i=1}^{n}\frac{1}{m}[D_{i}]$ and [24, Theorem 1.7] implies that the Fujita invariant is $n/m$, while the $b$-invariant is equal to